State the ordered pair that satisfies the equation
2x-5y=-1

Answer by rapaljer(4671)

State "the ordered pair"?? There are infinitely many ordered pairs that satify this equation. Perhaps you are just looking for one such pair? If so, then pick y = 1. Then 2x -5=-1, so 2x = 4, and x= 2.

How about that: (2, 1).

Check:
2x-5y = -1
2*2 -5*1 =-1
4-5=-1

It checks!!

R^2